OpenStudy (anonymous):

Parallel lines are cut by a transversal such that the alternate interior angles have measures of 3x + 17 and x + 53 degrees. Find the value of x.

OpenStudy (imstuck):

Do you have the image for this?

OpenStudy (imstuck):

If they are what I think they are drawn in this way:then 3x+17=x+53

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

Alternate interior angles, if a transversal cuts 2 parallels are equal to each other... So, 3x + 17 = x + 53

OpenStudy (anonymous):

Solve that for x.

OpenStudy (anonymous):

the options are 9,18,35,71

OpenStudy (imstuck):

See, like I said from my drawing, alternate interior angles are equivalent. So 3x+17=x+53. Solve that for x. Can you do that?

OpenStudy (anonymous):

can you just tell me the fluttering answer